### v1

# Training Large Language Models to Reason In Parallel with Global Forking Tokens

Abstract
Although LLMs have demonstrated improved performance by scaling parallel test-time compute, doing so relies on generating reasoning paths that are both diverse and accurate. For challenging problems, the forking tokens that trigger diverse yet correct reasoning modes are typically deep in the sampling tree. Consequently, common strategies to encourage diversity, such as temperature scaling, encounter a worsened trade-off between diversity and accuracy. Motivated by this challenge, we treat parallel reasoning as a set-of-next-token-prediction problem, and incorporate a set-based global loss into Supervised Fine-Tuning (SFT) using self-supervised bipartite matching between our global forking tokens and unique reasoning traces. We observe that, while naive fine-tuning with multiple reasoning traces collapses these unique reasoning modes, our proposed method, Set Supervised Fine-Tuning (SSFT), preserves these modes and produces emergent global forking tokens. Experiments on multiple reasoning benchmarks show that our SSFT consistently outperforms SFT under both $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ metrics.
Figure 1: An illustration of different supervised fine-tuning methods that aim to instill parallel reasoning capabilities from diverse reasoning traces. Compared to (1) standard SFT and (2) SFT with randomly assigned parallel thinking identifiers, (3) Set-Supervised Fine-Tuning leverages a self-supervised bipartite matching process to learn to initiate distinct reasoning modes conditioned on distinct <think i>, for each question. Self-supervised matching prevents collapse of distinct reasoning modes caused by ordering bias in a set of parallel reasoning traces.
1 Introduction
Large language models have recently improved reasoning by allocating more test-time compute to generate more tokens before producing the final answer (OpenAI, 2025). However, extended sequential scaling can lead to “overthinking”, where performance decreases beyond a certain sequence length (Ghosal et al., 2025; Chen et al., 2024a). To mitigate this, another scaling dimension based on repeated parallel sampling and aggregation (Wang et al., 2022; Brown et al., 2024) has shown success in further boosting reasoning performance. However, these methods rely on LLMs generating diverse yet correct solutions; as tasks become harder, a new mechanism is required to increase diversity. Recent work shows that only a minority of tokens in Chain-of-Thought reasoning (Wei et al., 2022) can act as forking tokens that lead to distinct reasoning modes (Wang et al., 2025), so as the problem becomes harder and the generation becomes longer, it can become substantially harder to sample them. Also, common practices to encourage diversity, typically through temperature scaling, inherently entail a diversity-accuracy trade-off, as the forking tokens that trigger diverse yet correct reasoning modes are typically located deeply within the sampling tree. Moreover, recent theoretical work also shows that increasing the temperature alone does not necessarily guarantee greater diversity unless the model is explicitly trained to ensure coverage. (Verine et al., 2025).
Building on these observations, we aim to leverage diverse reasoning traces to train for coverage (Guo et al., 2025; Google, 2025b). We introduce global forking tokens to initiate parallel reasoning traces and frame parallel reasoning as a set prediction problem. Specifically, given a question, an LLM, conditioned on a reserved set of tokens in a chosen ordering, generates $\mathrm{M}$ reasoning sequences in parallel, each aligned with one of $\mathrm{M}$ ground-truth reasoning traces. For each ordering, we compute the total autoregressive loss across the generated sequences. By enumerating all possible orderings, we identify the minimum loss, which defines the set language modeling loss, conditioned on the distinct forking tokens (Equation 5). This formulation naturally incorporates coverage into the training objective and is capable of learning global forking tokens that can serve as prompts to trigger reasoning modes that are both diverse and accurate. We operationalize this idea through our Set Supervised Fine-Tuning (SSFT) framework. Our main contributions are summarized below. We plan to open-source our code and model.
  • We introduce global forking tokens and incorporate a set-based loss into SFT using bipartite matching between a set of reserved global forking tokens and diverse reasoning traces (Section 2; Figure 2). The goal is that, after training with SSFT, the model can initiate distinct reasoning modes when prompted with different tokens from this set, thus reducing the dependence on sampling forking tokens mid-generation (Section 2.3; Figure 3).
  • We empirically demonstrate that, across multiple reasoning benchmarks, a model fine-tuned with SSFT outperforms a standard SFT model trained on the same distilled traces from diverse teacher models, improving $\mathrm{Pass@}1$, $\mathrm{Pass@}k$, and $\mathrm{Cons@}k$ (Section 3, Table 1 and Figure 4). We also find that naively fine-tuning using diverse reasoning traces without the set loss from bipartite matching can result in control tokens initiating the same collapsed reasoning mode (Figure 6). In contrast, the global forking tokens learned by SSFT reliably initiate distinct reasoning (Figure 5).
  • To facilitate training with variable-size parallel generation, we propose a scalable training implementation that does not increase VRAM usage. Instead of concatenating diverse reasoning traces, our algorithm can expand these variable-sized parallel generations along the batch dimension under distributed training (Appendix A.1).
2 Learning Global Forking Tokens via Set Supervised Finetuning
Background on Language Modeling with Reasoning. In language modeling, the goal is to train a model $\pi_{\bm{\theta}}$ to approximate the joint distribution over a sequence of word tokens ${\mathbf{x}}=\{x_{i}\}_{i=1}^{\mathrm{T}}\in\mathcal{V}^{\mathrm{T}}$, where $\mathrm{T}$ is the sequence length, and each token is within a finite vocabulary set $\mathcal{V}$. An autoregressive model uses the chain rule to represent it as a product of conditionals on the preceding tokens:
$\pi_{\bm{\theta}}({\mathbf{x}})=\prod_{t=1}^{\mathrm{T}}\pi_{\bm{\theta}}\left({\textnormal{x}}_{t}|{\mathbf{x}}_{<t}\right)$ (1)
[…]
2.1 Parallel Reasoning as Set of Next Token Prediction
Problem Setup. In this paper, our goal is not only to instill new reasoning capabilities into a model, but also to ensure that prompting with a set of reserved special tokens, in parallel with a question, elicits distinct reasoning traces. We call these global forking tokens ${\bm{g}}:=\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ instantiated as $\{\texttt{<think\,i>\,}\}_{i=1}^{\mathrm{N}}$ tags. We use ${\textnormal{g}}^{(i)}$ interchangeably with <thinki>, depending on context for clarity. We consider a setting with multiple sources of reasoning traces, obtained at low cost without human annotation by distilling from diverse teachers, sampling repeatedly, and potentially filtering with a verifiable metric such as correctness. We adopt this low-cost regime to highlight the effectiveness of our algorithm, though the method extends naturally to settings with well-annotated, human-labeled data. So our problem is to (1) learn to do a set of next-token predictions on multiple distinct yet correct reasoning traces ${\mathbf{R}}:=\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ in parallel for an input prompt $\mathbf{x}$}, and (2) ensure that distinct global forking tokens can uniquely initiate these distinct traces.
To do this, we make a simple change to the NTP loss, which now has two requirements: (1) Permutation-invariance: It should not depend on the order of elements in $\mathbf{R}$} and $\bm{g}$}, so we don’t penalize a trace that incurs high NTP loss under one forking token if the model predicts it well when conditioned on another. (2)No shared global forking token: We want $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ to uniquely initiate distinct reasoning traces, so this requirement prevents conditioning on the same ${\textnormal{g}}^{(i)}$ when generating distinct traces given a question.
To satisfy these requirements, we incorporate a subproblem in language modeling: finding the minimum cost bipartite matching configuration between the left vertices $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and the right vertices $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ where the cost of each edge between a left vertex $i$ and a right vertex $j$, is the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ and an input prompt $\mathbf{x}$}. A matching configuration is a set of edges connecting the left and right vertices where no two edges share a common vertex. Without loss of generality, we assume $\mathrm{N}\geq\mathrm{M}$ to simplify our notation. The total cost involves all vertices on the smaller side of the bipartite graph, so this allows us to write the summation from 1 to $\min\{\mathrm{N},\mathrm{M}\}$, which equals $\mathrm{M}$ under this assumption. We denote a matching configuration as a finite map $\bm{\sigma}:\{1,...,\mathrm{M}\}\rightarrow\{1,...,\mathrm{N}\}$ such that $\bm{\sigma}(j)=i\iff{\mathbf{r}}^{(j)}\text{ is paired with }{\textnormal{g}}^{(i)}$. Let $\bm{\mathfrak{S}}_{\mathrm{P}}:=\{\bm{\sigma}_{k}\}_{k=1}^{\mathrm{P}}$ denote all the $\binom{\mathrm{N}}{\mathrm{M}}\times\mathrm{M}!$ configurations of a bipartite graph. The total cost of each configuration represents the compatibility between $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ and $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ under this unique matching. Figure 2 shows an example of a matching configuration.
Figure 2: An illustration of one SSFT training step. Step 1: We first construct the cost matrix by evaluating all pairwise combinations: for each ${\mathbf{r}}^{(j)}\in\{{\mathbf{r}}^{(1)},{\mathbf{r}}^{(2)},{\mathbf{r}}^{(3)},{\mathbf{r}}^{(4)}\}$ and each ${\textnormal{g}}^{(i)}\in\{{\textnormal{g}}^{(1)},{\textnormal{g}}^{(2)},{\textnormal{g}}^{(3)},{\textnormal{g}}^{(4)},{\textnormal{g}}^{(5)},{\textnormal{g}}^{(6)}\}$, we compute the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ (Equation (4)). Then we use Hungarian algorithm to find $\hat{\bm{\sigma}}$ that minimizes the total bipartite matching cost. Here, this minimum is the sum of the losses highlighted in blue, which means $\hat{\bm{\sigma}}=\{({\textnormal{g}}^{(6)},{\mathbf{r}}^{(1)}),({\textnormal{g}}^{(5)},{\mathbf{r}}^{(2)}),({\textnormal{g}}^{(2)},{\mathbf{r}}^{(3)}),({\textnormal{g}}^{(3)},{\mathbf{r}}^{(4)})\}$. Step 2: We optimize $\bm{\theta}$} by backpropagating the set of NTP losses for ${\mathbf{r}}^{(j)}$, each conditioned on ${\textnormal{g}}^{(\hat{\bm{\sigma}}(j))}$.
2.2 SSFT: Minimizing Set of NTP Losses under Optimal Bipartite Matching
Under this formulation, we propose Set Supervised Fine-Tuning (SSFT), which performs two operations at each training step: (1) find the minimum-cost bipartite matching that is optimal for uniquely initiating different reasoning traces (2) and then minimize the NTP losses under the matching configuration to instill diverse reasoning modes conditioned on the matched global forking tokens. We show our implementation in Algorithm 1.
For the first step, we first compute all the entries in the cost matrix such as the one in Figure 2 and then apply the Hungarian algorithm (Kuhn, 1955) to efficiently find the optimal $\hat{\bm{\sigma}}$:
$\hat{\bm{\sigma}}=\operatorname*{arg\,min}_{\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}}\sum_{j=1}^{\mathrm{M}}\mathcal{L}_{\text{matching}}\!\left({\textnormal{g}}^{(\bm{\sigma}(j))},{\mathbf{r}}^{(j)}\right),\,\mbox{ where }$ (3)
$\mathcal{L}_{\text{matching}}\!\left({\textnormal{g}}^{(i)},{\mathbf{r}}^{(j)}\right)=-\text{sg}\left(\frac{1}{\mathrm{T}_{\mathbf{r}}}\sum_{t=1}^{\mathrm{T}_{{\mathbf{r}}}}\log\pi_{\bm{\theta}}\left({\textnormal{r}}_{t}^{(j)}|{\mathbf{x}},{\textnormal{g}}^{(i)},{\mathbf{r}}^{(j)}_{<t}\right)\right)$ (4)
As noted in Equation 4, each matching cost in Equation 3 is the negative log-likelihood of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(\bm{\sigma}(j))}$ under the current model parameters. Here, $\text{sg}(\cdot)$ is stop-gradient, as the matching process is done by discrete optimization w.r.t. $\bm{\sigma}$, so we can save VRAM by not storing intermediate activations. We explicitly indicate that length normalization is done to remove biases toward trace length, so that the matching is driven by semantic content.
After solving $\hat{\bm{\sigma}}$ for each $({\mathbf{x}},{\mathbf{R}})$, our second step optimizes model parameters $\bm{\theta}$} by backpropagating on the matching loss in Equation 5. The expectation is replaced by its sample mean over pairs of $({\mathbf{x}},\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}})$ in a mini-batch. In practice, we may use only the first $\mathrm{L}<{\mathrm{T}_{{\mathbf{r}}}}$ tokens in Equation 4 to compute the matching cost and find $\hat{\bm{\sigma}}$ when the training dataset is such that these early tokens reveal sufficient differences. However, we always optimize the matching loss for the full $\mathrm{T}_{{\mathbf{r}}}$} length.
$\mathcal{L}_{\text{Hungarian}}\!\left({\bm{\theta}}\right)=-\operatorname*{\mathbb{E}}_{{\mathbf{x}},{\mathbf{R}}\sim\mathcal{D}}\left[\sum_{j=1}^{\mathrm{M}}\sum_{t=1}^{{\mathrm{T}_{{\mathbf{r}}}}}\log\pi_{\bm{\theta}}\left({\textnormal{r}}_{t}^{(j)}|{\mathbf{x}},{\textnormal{g}}^{(\hat{\bm{\sigma}}(j))},{\mathbf{r}}^{(j)}_{<t}\right)\right$] (5)
Remarks. The resulting model is not the same as a simple routing of the models independently trained with the nonoverlapping subsets of these traces. Firstly, SSFT allows positive transfer in representation learning within $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ even though they are matched to different $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$. Secondly, it is not optimal to distill reasoning traces from the same fixed sources for every question if the goal is to maximize both diversity and correctness. Our algorithm supports a variable number of target reasoning traces across training steps, with sources that may also change for each $\mathbf{x}$}. Thirdly, even if the two sets of traces, $\{{\mathbf{r}}^{(j)}_{a}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{a}$ and $\{{\mathbf{r}}^{(j)}_{b}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{b}$, are from the same sources, their optimal configurations $\hat{\bm{\sigma}}_{a}$ and $\hat{\bm{\sigma}}_{b}$ can still vary because a teacher model can reason differently under different questions. Lastly, we reserve more global forking tokens than the maximum number of traces ($\mathrm{N}>\mathrm{M}$), and empirically observe that all the forking tokens are being matched throughout the process. This is because the extra forking tokens can maximally intra-differentiate similar traces.
2.3 Inference with Learned Global Forking Tokens
We discuss the inference protocols with $\mathrm{N}$ global forking tokens.
Figure 3: (Left) $\bm{\mathfrak{S}}_{p}$ is the subset of bipartite matching configurations that are still computed optimal towards the end of training. These can be individually visualized as a bipartite (Middle) Learned matchings by SSFT-32B in Exp 3, obtained by aggregating all edges in $\bm{\mathfrak{S}}_{p}$. (Right) At test time, for $\mathrm{Pass@}1$, we prompt with ${\textnormal{g}}^{(i^{\star})}$ that has the most connected edges. For $\mathrm{Cons@}k$, we augment $i$th prompt by <think(i % $\mathrm{N}$)>.
[…]
$i^{\star}=\operatorname*{arg\,max}_{i}\lvert\bigcup_{\bm{\sigma}\in\bm{\mathfrak{S}}_{p}}{\{j|\bm{\sigma}(j)=i\}}\rvert$ (6)
3 Experiments
We address the following research questions through experiments: (RQ1): In terms of $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ accuracy, how does a model trained with SSFT perform on reasoning benchmarks? (RQ2):  Does finding the optimal bipartite matching matter in reasoning performance? (RQ3): Does training with diverse reasoning traces yield better accuracy and coverage under SSFT compared to standard SFT with temperature scaling? (RQ4): Does prompting with distinct $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ genuinely make a model generate diverse reasoning traces? (RQ5): Is the performance gain from SSFT conditioned on the traces generated by our procedure and on a high-quality small dataset?
3.1 Experiment Setup
Training Dataset. We use the 1,000 questions from s1k dataset Muennighoff et al. (2025). In addition to the R1 (Guo et al., 2025) and Gemini Flash (Google, 2024) traces provided by s1, we also use Claude Opus 4.0/4.1 (Anthropic, 2025) and GPT-OSS-120B (Agarwal et al., 2025) with high and medium reasoning effort to obtain a pool of distilled targets for the 1,000 questions. For each question, we generate two traces per source to populate the pool. We then sample four traces from this pool. We call this s1k-4mixed-reasoning dataset.
Training Details. We fine-tune Qwen2.5-32B-Instruct (Yang et al., 2025a) for six epochs with a context length of 32,768. We reserve $\mathrm{N}=6$ global forking tokens and use $\mathrm{M}=4$ targets per question. To find the optimal bipartite matching for each input prompt, we consider only the first 1,000 tokens when computing the matching cost in Equation 4 for computational efficiency. We call this model SSFT-32B. We also include SSFT but choose a random bipartite matching at each step to fine-tune SSFT-32B (random $\bm{\sigma}$). Exact details on the pool of diverse distillation targets and selection procedure, as well as training hyperparameters, are provided in Appendix A.3.
Baselines. All of our baselines use Qwen2.5-32B-Instruct as their base model, and only train on the 1k questions. Our baselines include two groups: (Single-Target $\triangle$) models trained with one trace per question and (Multi-Target $\bigstar$) models trained with four traces per question. For (Single-Target), we include s1.1-32B (Muennighoff et al., 2025), which uses 1k DeepSeek-R1 traces. We also fine-tuned an SFT-OSS-distill-32B baseline that trains only on the 1k GPT-OSS traces with high reasoning effort, as these traces achieved the highest correctness on the 1k questions based on an evaluation by Claude 3.5 Sonnet comparing each attempt against the reference answer. For (Multi-Target), we use our s1k-4mixed-reasoning to fine-tune SFT-mixed-distill-32B using standard SFT with one <think> token, duplicating each question and treating the four traces as four individual data points. We also include Multiverse-32B (Yang et al., 2025b), which prompts Gemini 2.5-Pro (Google, 2025a) to transform 1k sequential CoTs into parallel CoTs as their training data.
Evaluation Setup. Our evaluation tasks consist of AIME24/AIME25 (Ye et al., 2025), MATH-500 (Hendrycks et al., 2021), and GPQA-Diamond (Rein et al., 2024). We use LightEval (Habib et al., 2023) as our evaluation framework with generation configurations: temperature=0.7 used in (Guha et al., 2025), top_p=0.95, max length=32768. For $\mathrm{Pass@}1$ accuracy without any parallel test-time compute, we select learned ${\textnormal{g}}^{(1)}$ for SSFT-32B and ${\textnormal{g}}^{(4)}$ for SSFT-32B (random $\bm{\sigma}$) based on Equation 6. For each $\mathrm{Pass@}1$ accuracy, we compute the average performance over 32 generations. For $\mathrm{Cons@}6$, which applies each of the six global forking tokens once in our method and uses six generations for the baselines, we compute the average over 11 sets of generations to reduce variance in the results. We refer to this as Pass@1 of Native Cons@6 using a similar terminology as the concurrent work (Wen et al., 2025). Appendix  A.4 presents an example of the parallel generations.
3.2 Evaluating SSFT on Reasoning Benchmarks
AIME 2024 AIME 2025 MATH-500 GPQA-D Average
Pass@1: Average performance of individual generations
Qwen2.5-32B-Instruct $\triangle$ 15.80 10.40 80.40 47.00 38.40
s1.1-32B $\triangle$ 54.79 44.27 92.16 62.12 63.34
Multiverse-32B $\bigstar$ 53.80 45.80 91.80 60.70 63.03
[…]
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 61.77 55.10 89.95 62.28 67.28
SSFT-32B $\bigstar$ 64.06 58.13 90.02 60.39 68.15
Pass@1 of Native Cons@6: Average performance of majority voting with 6 parallel generations
s1.1-32B $\triangle$ 70.30 53.33 95.60 61.45 70.17
[…]
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 73.03 67.58 95.67 61.87 74.54
SSFT-32B $\bigstar$ 75.45 73.94 96.47 63.05 77.23
Cons@32: Majority voting performance with large number of parallel generations
s1.1-32B $\triangle$ 73.33 63.33 94.80 60.61 73.02
[…]
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 80.00 80.00 95.60 62.63 79.56
SSFT-32B $\bigstar$ 83.33 86.67 96.80 61.62 82.11
$\triangle$ indicates training with single-target data and $\bigstar$ indicates training with multi-target data.
Table 1: Performance of SSFT compared to baselines on four reasoning tasks, reported at $\mathrm{Pass@}1$, $\mathrm{Cons@}6$, and $\mathrm{Cons@}32$. SSFT selects <think1> for $\mathrm{Pass@}1$ and replaces 6 generations with 6 generations prompted by distinct <thinki> for $\mathrm{Cons@}k$. We observe consistent improvements over (i) SFT-OSS-distill-32B, which uses the 1k OSS-high traces; (ii) SFT-mixed-distill-32B, which uses the four mixed traces but treats them as individual data; and (iii) SSFT-32B (random $\bm{\sigma}$), which trains using Equation 5 but with a randomly chosen $\bm{\sigma}$.
For RQ1, we see in Table 1 that SSFT delivers the best $\mathrm{Pass@}1$ accuracy, 64.06 on AIME24 and 58.13 on AIME25, outperforming SFT-mixed-32B by $8.33\%$ and $6.57\%$, respectively. We also observe consistent improvements on all four tasks under parallel test-time compute at two scales, $\mathrm{Cons@}6$ and $\mathrm{Cons@}32$, over SFT-mixed-32B, which was trained on the same reasoning traces. Some notable results are $\mathrm{Cons@}6=73.94\%,\mathrm{Cons@}32=86.67\%$ on AIME25. To answer RQ2, we observe consistent improvements over SSFT-32B (random $\bm{\sigma}$), with especially strong gains at $\mathrm{Cons@}6$ on AIME25, where effectiveness with few parallel generations is critical. As shown later in Figures 5 and  6, optimal bipartite matching is essential for preventing collapsing reasoning modes.
(a) AIME24
(b) AIME25
Figure 4: Coverage of SSFT compared to SFT-mixed-distill-32B with temperature scaling, reported at $\mathrm{Pass@}k$. For convenience, we also report the $\mathrm{Cons@}6$ accuracy next to each line. In AIME25, SFT-mixed-distill-32B needs to raise the inference temperature to 1 and use more attempts to match the coverage at the cost of lowering its $\mathrm{Pass@}1$ and $\mathrm{Cons@}6$ accuracy, further widening the gaps.
For RQ3, we compare our method against SFT-mixed-32B under various $k$ in $\mathrm{Pass@}k$ accuracy with 32 generations to assess generation coverage. Figure 4 shows that SSFT achieves higher coverage across nearly all values of $k$. SFT-mixed-32B requires more allowed attempts and higher temperature to match the coverage of SSFT at the cost of lowering its $\mathrm{Pass@}1$ and $\mathrm{Cons@}6$ accuracy.
3.3 Evaluating Parallel Reasoning Diversity and Learned Matchings
Figure 5: (SSFT, optimal matching). Distribution of thinking-token counts and average performance on AIME24 (left) and AIME25 (right) prompted by a distinct $\texttt{<think1>},\dots,\texttt{<think6>}$. Each accuracy is averaged over 11 generations, shown above each whisker. We observe clear specialization of reasoning modes. $\texttt{<think1>},\dots,\texttt{<think4>}$ trigger longer traces, better suited to many AIME problems, and these surpass all the average accuracy under SFT without bipartite matching in Figure 6. <think5> and <think6> favor concise reasoning, which can be suboptimal for some AIME tasks, yet still raise $\mathrm{Cons@}6$ beyond all baselines by adding diversity. The consistency of these length distributions across AIME24 and AIME25 indicates the difference is not due to randomness, and that these think tags truly initiate distinct yet consistent reasoning modes.
Figure 6: (SSFT, random matching). Distribution of thinking-token counts and average performance on AIME24 (left) and AIME25 (right) prompted by a distinct $\texttt{<think1>},\dots,\texttt{<think6>}$. Each accuracy is averaged over 11 generations, shown above each whisker. Without optimal bipartite matching during SFT, the think tags initiate reasoning traces with similar length distributions and achieve similar average performance. Moreover, the slight length differences are not consistent across AIME24 and AIME25, suggesting randomness rather than genuine specialization; the tags therefore collapse to similar modes, failing to preserve the diverse reasoning modes in the training data. Both the average performance and $\mathrm{Cons@}6$ fall behind those from any think tag that initiates long reasoning under SSFT in Figure 5.
Addressing RQ4, we show that our global forking tokens genuinely initiate distinct reasoning traces and offer a new mechanism for leveraging test-time compute.
Emerging Diverse Reasoning Modes. Using the $\mathrm{Cons@}6$ results in Table 1, we form six sets of generations, each prompted by a distinct ${\textnormal{g}}^{(i)}$. For each set, we show the average accuracy and the distribution of thinking-token counts: Figure 5 for SSFT with optimal bipartite matching and Figure 6 for SSFT with random matching. Reasoning length: Length partially indicates the diversity in reasoning, and we see clear differences for SSFT with optimal matching, despite the absence of hand-crafted matching rule or information about these traces. The consistency of these distributions across AIME24 and AIME25 indicates the differences is not from randomness, whereas randomly assigning a <thinki>, as in concurrent work (Wen et al., 2025), does not yield clear or consistent differences in reasoning length, as shown in Figure 6. Performance: After finetuning with random matching, prompting with a distinct <thinki> shows no meaningful impact ($\approx 61\%$ on AIME24 and $\approx 55\%$ on AIME25). With optimal matching, SSFT elicits distinct reasoning modes initiated by $\texttt{<think1>},\dots,\texttt{<think4>}$ that reach around $65\%$ on AIME24 and $\geq 59\%$ on AIME25, with different lengths. Although <think5> and <think6> are weaker due to shorter reasoning modes, the average of these and, especially, $\mathrm{Cons@}6$ performance improve consistently with them.
(a) SSFT under random matching ($\{{\textnormal{g}}^{(i)}\}_{i=1}^{4},\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$). No correlations learned $\bm{\mathfrak{S}}_{p}=\bm{\mathfrak{S}}_{\mathrm{P}}$
(b) SSFT under optimal matching ($\{{\textnormal{g}}^{(i)}\}_{i=1}^{4},\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$). $\bm{\mathfrak{S}}_{p}=\{22,23\}$
(c) SSFT under optimal matching ($\{{\textnormal{g}}^{(i)}\}_{i=1}^{6},\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$). $\bm{\mathfrak{S}}_{p}=\{62,110,134,180,204,230\}$
Figure 7: Cumulative counts of $\bm{\sigma}_{k}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ computed as optimal over training. Note that $\bm{\mathfrak{S}}_{\mathrm{P}}$ and $\bm{\mathfrak{S}}_{p}$ are defined in Sections 2.1 and 2.3, respectively. Front axis: matching configuration index $k$. Depth: training step $t$. Bar height is the cumulative counts. These are the evolution of matchings during training 3 Qwen-32B-Instruct models under 3 bipartite matching settings. In this case study, the $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$ are always (GPT-OSS-high, GPT-OSS-med, R1, Gemini) for each question. Note that Random matching method does not minimize Eqn 5 under optimal matching, but we track it. We observe $\bm{\mathfrak{S}}_{p}=\bm{\mathfrak{S}}_{\mathrm{P}}$ with random matching, meaning no correlations learned. But by optimizing Hungarian loss, we see the emergence of $\bm{\mathfrak{S}}_{p}\subset\bm{\mathfrak{S}}_{\mathrm{P}}$.
(d) Visualization of learned matchings after SSFT with random matching. The fully connected graph indicates no correlations were learned.
(e) Visualization of learned matchings after SSFT with optimal matching between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
(f) Visualization of learned matchings after SSFT with optimal matching between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
Figure 8: We visualize the learned matchings by connecting edges in the subset of configurations $\bm{\mathfrak{S}}_{p}$ that still accumulate mass late in training (Figure 7). All the three models in (a), (b), and (c) are fine-tuned using GPT-OSS-high, GPT-OSS-medium, R1, and Gemini to distill reasoning traces for every question, enabling clearer interpretation.
Visualization of Learned Matchings. We observe in Figure 3 that some ${\textnormal{g}}^{(i)}\in{\bm{g}}$ from SSFT-32B have unique configuration of matched edges with $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. This is a positive indication that $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ are likely to initiate distinct reasoning modes. We hypothesize that $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ can still yield a unique edge-matching configuration even if a subset of $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$ are difficult to distinguish.
For interpretability, we fine-tune using the same four teacher models for each $\mathbf{x}$}: GPT-OSS high, medium, R1, and Gemini. We call this dataset s1k-4teachers-reasoning dataset, and ask whether SSFT associate a unique ${\textnormal{g}}^{(i)}$ to the Gemini and R1 traces which have easily identifiable reasoning patterns, and then have the rest matched to OSS-high/OSS-medium traces in different ways (i.e. only matched to OSS-high, only matched to OSS-med, and matched to both). We track source indices only for evaluation; the model still receives an unordered set of traces with no source information for each question. We study three bipartite matching settings with SSFT. Figure 7 shows the evolution of matchings learned under these three hyperparameters: (a) random matching with four reserved ${\textnormal{g}}^{(i)}$, (b) optimal matching with four reserved ${\textnormal{g}}^{(i)}$, and (c) optimal matching with six reserved ${\textnormal{g}}^{(i)}$. Initially, all the configurations $\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ accumulate mass as there is no correlations between $\bm{g}$} sand $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. Figure 7(a) shows SSFT under random matching does not shrink the size of configurations computed as optimal, meaning that no correlations are learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. By contrast, Figure 7(b) and Figure 7(c) show the emergence of only a strict subset of matching configurations in $\bm{\mathfrak{S}}_{\mathrm{P}}$. This indicates some correlations between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ are indeed learned through SSFT. We first visualize the learned matchings of the model with four ${\textnormal{g}}^{(i)}$ in Figure 8. We observe that ${\textnormal{g}}^{(1)}$ and ${\textnormal{g}}^{(2)}$ are uniquely matched to the R1 and Gemini traces, showing that SSFT can indeed uniquely associate ${\textnormal{g}}^{(i)}$ to sufficiently diverse reasoning traces. Now to confirm our previous hypothesis, we see the unique learned matchings $({\textnormal{g}}^{(3)},(\text{OSS-high, OSS-med})),({\textnormal{g}}^{(4)},\text{OSS-med})$. Furthermore, by connecting all the edges in $\bm{\mathfrak{S}}_{p}$ from SSFT with 6 forking tokens (Figure 7(c)), we also see unique learned matchings $({\textnormal{g}}^{(3)},\text{OSS-high}),({\textnormal{g}}^{(4)},\text{OSS-med}),({\textnormal{g}}^{(5)},(\text{OSS-med},\text{OSS-high}))$ in Figure 8. This confirms that the global forking tokens can identify unique correlations even among highly similar traces.
3.4 Ablation Study: Removing High Quality Small Dataset
To test whether our empirical gains are conditioned on the traces generated by our procedure and on a highly optimized small dataset (Muennighoff et al., 2025), we fine-tune on a public dataset that already provides sufficient reasoning traces per question: the 93k math set of Face (2025) (2-4 traces per question). Because this dataset has been successful for fine-tuning Qwen2.5-Math-7B, we adopt that base model and compare SSFT against SFT trained on all available traces. Details on the hyperparameters are in Appendix A.3.4. Addressing RQ5, Table 2 shows consistent improvements in both $\mathrm{Pass@}1$ and $\mathrm{Cons@}32$. The results indicate the SSFT is effective for larger sized public dataset with less diverse reasoning traces.
AIME 2024 AIME 2025 MATH-500 GPQA-D
Model Pass@1 Cons@32 Pass@1 Cons@32 Pass@1 Cons@32 Pass@1 Cons@32
Qwen2.5-Math-7B-Instruct 10.42 20.00 9.48 23.33 81.87 87.40 30.29 30.30
SFT-OpenR1-93k-7B 46.15 66.67 34.17 50.00 86.62 90.20 46.35 47.98
SSFT-OpenR1-93k-7B 51.25 73.33 35.52 56.66 89.74 93.60 46.86 48.90
Table 2: Performance of SSFT versus SFT trained solely on publicly available distillation targets. The setup uses the 93k math questions from Face (2025) with Qwen2.5-Math-7B as the base model. SFT-OpenR1-93k-7B uses the same distillation targets as SFT-mixed-distill-32B in Table 1.
4 Related work
Test-time Scaling. There has been a surge of work fine-tuning LLMs to reason longer, using reinforcement learning for frontier models (OpenAI, 2024; Shao et al., 2024; xAI, 2025; Yang et al., 2025a) and supervised fine-tuning for smaller ones (Muennighoff et al., 2025; Hu et al., 2025). These methods enable LLMs to improve reasoning by allocating more test-time compute to sequential, iterative refinement such as self-reflection (Guo et al., 2025; Liu et al., 2025). However, extended sequential reasoning can be more sensitive to the order of reasoning steps and may result in failures (Chen et al., 2024b), and performance can start to degrade beyond a certain length due to “overthinking” (Ghosal et al., 2025). Our goal is to study the effective use of diverse reasoning traces to fine-tune small language models, essential for agentic AI (Belcak et al., 2025).
Parallel Reasoning. Parallel scaling methods such as self-consistency  (Wang et al., 2022) and Best-of-N (Lightman et al., 2023) improve LLM performance by generating multiple reasoning paths in parallel and aggregating them. These methods fundamentally require choosing a temperature that can generate diverse reasoning paths, but a recent theoretical work shows that increasing temperature can sometimes fail to increase diversity if language models are not trained towards coverage (Verine et al., 2025). Other search-based methods such as Monte Carlo tree search (MCTS) (Zhang et al., 2024) and Tree of Thoughts (ToT) (Yao et al., 2023) apply heuristic-guided search with an external verifier to do more deliberate search to increase the coverage. However, their dependence on heuristics and domain-specific knowledge can limit their applicable tasks. Regarding training LLMs with parallel reasoning traces,  Yang et al. (2025b) proposes training with parallel CoTs decomposed from sequential CoTs, and our concurrent work Wen et al. (2025) proposes to train with multiple reasoning traces distilled from teacher models. These works show native parallel scaling can surpass sequential scaling within certain token limits. Although we also show improvements in reasoning performance by simply fine-tuning with parallel reasoning traces, we empirically highlighted the limitation of manually assigning parallel think tags to target reasoning traces. If human assignment or random enforcement routes mismatched traces to a tag with a unique reasoning mode learned, the tag’s semantics blur due to bias in the ordering of parallel reasoning traces. This prevents tags from initiating distinct reasoning modes. We are the only work that leverages parallel think tags to construct a set-prediction loss via bipartite-matching optimization between tags and reasoning traces, turning the tags into global forking tokens. Empirically, we also improve $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ over baselines fine-tuned with parallel reasoning traces on the same dataset, whether on subsets or the full set. Another recent work (Zheng et al., 2025) encourages the use of parallel thinking tags by slightly upweighting rewards for correct trajectories that employ such tags during reinforcement learning. However, the cold-start SFT that instills parallel thinking from parallel reasoning traces is still performed without a set-prediction loss; as a result, the tags may initiate similar reasoning traces.
Set-based Global Loss in Deep Learning. DETR introduces end-to-end object detection with a set global loss (Carion et al., 2020; Minderer et al., 2022), whose success in parallel bounding-box prediction inspires our approach. We are the first to extend this to language modeling: While DETR predicts a set of single tokens in parallel to match a list of bounding boxes, we predict a set of sequential thinking tokens, initiated by global forking tokens, and evaluate the matchings using autoregressive losses. We adapt set-based loss with autoregressive (AR) models, rather than diffusion-based models (Zhao et al., 2025; Nie et al., 2025; Ho et al., 2020), because AR models achieve superior reasoning performance. Several innovations are needed to make this loss practical for parallel reasoning with AR models. For example, to populate the cost matrix we need $\mathrm{N}\times\mathrm{M}$ forward passes, one for each combination of $\mathrm{N}$ think tags and $\mathrm{M}$ reasoning traces, whereas DETR requires only $\mathrm{N}$ forward passes for its $\mathrm{N}$ queries, since the same bounding-box predictions can be reused to compute matching costs against different ground-truth boxes. We therefore emphasize efficient training implementation in our extension, increasing training time by only $6.6\%$ compared to standard SFT (Appendix A.1).
5 Conclusion
In this work, we demonstrate that diverse reasoning traces can be leveraged to learn global forking tokens which serve as prompts to initiate distinct reasoning modes that are both diverse and accurate. We propose Set Supervised Fine-Tuning (SSFT), which employs bipartite matching between reserved global forking tokens and diverse reasoning traces to compute a set-based language modeling loss. We show that models trained with SSFT yield improvements in both $\mathrm{Pass@}k$ and $\mathrm{Cons@}k$ accuracy over various $k$ compared to standard SFT with temperature scaling. The proposed training method also improves the usage of multi-target data and avoids collapsing distinct reasoning modes, which is especially helpful in the context of distilling from multiple teacher models. In future work, we plan to further investigate how performance scales with a larger number of distillation targets and reserved forking tokens, which jointly determine the size of the underlying bipartite graph. In addition, we will apply RFT to a model pre-trained with SSFT to incentivize optimal use of the learned global forking tokens across a broader suite of tasks without requiring additional high-quality reasoning traces.
Reproducibility. To ensure reproducibility of our work, we provide detailed hyper-parameter settings in A.3.2, and SSFT algorithm implementations in A.1. We are preparing the upcoming open-source release of our code and model.
Appendix A Appendix
In this appendix, we provide details omitted from the main text.
A.1 Algorithm: SSFT Implementation
Algorithm 1 Set Supervised Fine-tuning (SSFT)
1: $\pi_{{\bm{\theta}}}$: base model $\mathrm{N}$: Number of global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ $\mathcal{D}$: Dataset with (at most) $\mathrm{M}$ reasoning traces per question $\mathrm{B}$: Global batch size $\mathrm{T}_{\mathrm{L}}$}: The first $\mathrm{T}_{\mathrm{L}}$} number of tokens to match in matching cost (Equation  7).
2:Output $\pi_{\bm{\theta}}$
3:for each training step do
4:  for $k=1,...,\mathrm{B}$ do
[…]
6:   Initialize cost matrix $\mathrm{C}\in\mathbb{R}^{\mathrm{N}\times\mathrm{M}}$
7:   for $i=1,...,\mathrm{N}$ do
8:     for $j=1,...,\mathrm{M}$ do
9:      Compute the matching cost between ${\textnormal{g}}^{(i)}$ and ${\mathbf{r}}^{(j)}_{k}$ by Equation 7.
$\mathcal{L}_{\text{matching}}\!\left({\textnormal{g}}^{(i)},{\mathbf{r}}^{(j)}_{k}\right)=-\text{sg}\left(\frac{1}{{\mathrm{T}_{\mathrm{L}}}}\sum_{t=1}^{\mathrm{T}_{\mathrm{L}}}\log\pi_{\bm{\theta}}\left({\textnormal{r}}_{k,t}^{(j)}|{\mathbf{x}}_{k},{\textnormal{g}}^{(i)},{\mathbf{r}}^{(j)}_{k,<t}\right)\right)$ (7)
10:      Store the matching cost in $\mathrm{C}$.
$\mathrm{C}(i,j)=\mathcal{L}_{\text{matching}}\!\left({\textnormal{g}}^{(i)},{\mathbf{r}}^{(j)}_{k}\right)$ (8)
11:     end for
12:   end for
13:   Compute optimal matching $\hat{\bm{\sigma}}_{k}$ between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}_{k}^{(j)}\}_{j=1}^{\mathrm{M}}$. Hungarian algorithm (Kuhn, 1955) can be applied to $\mathrm{C}$ to efficiently compute Equation 9 (Equation 3).
$\hat{\bm{\sigma}}_{k}=\operatorname*{arg\,min}_{\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}}\sum_{j=1}^{\mathrm{M}}\mathrm{C}(\bm{\sigma}(j),j)$ (9)
14:  end for
15:  Compute the empirical set language modeling loss (Equation 10):
$\mathcal{L}_{\text{Hungarian}}\!\left({\bm{\theta}}\right)=-\frac{1}{\mathrm{B}}\sum_{k=1}^{\mathrm{B}}\left[\sum_{j=1}^{\mathrm{M}}\sum_{t=1}^{{\mathrm{T}_{{\mathbf{r}}}}}\log\pi_{\bm{\theta}}\left({\textnormal{r}}_{k,t}^{(j)}|{\mathbf{x}}_{k},{\textnormal{g}}^{(\hat{\bm{\sigma}}_{k}(j))},{\mathbf{r}}^{(j)}_{k,<t}\right)\right$] (10)
16:  Update model parameters $\bm{\theta}$} using gradients $\nabla_{{\bm{\theta}}}\mathcal{L}_{\text{Hungarian}}\!\left({\bm{\theta}}\right)$
17:end for
Algorithm 2 Queue-based Distributed SSFT with variable number of traces for each question
1: $\pi_{{\bm{\theta}}}$: base model $\mathrm{N}$: Number of global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ $\mathcal{D}$: Dataset with (at most) $\mathrm{M}$ reasoning traces per question, variable $m$ number of traces per question $\mathrm{B}$: Global batch size $\mathrm{T}_{\mathrm{L}}$}: The first $\mathrm{T}_{\mathrm{L}}$} number of tokens to match in matching cost (Equation  7). $b$: Original per-device global batch size ($b\geq\mathrm{M}$), This is “micro batch size*original grad accumulation steps”
2:Output $\pi_{\bm{\theta}}$
3:for each epoch do
[…]
5:  Initialize Queue $q$ for storing a sequence of sizes of sets $m$.
[…]
8:   $q\leftarrow q.enqueue(m)$
9:   $all\_q\_list=All\text{-}gather(q)$
10:   Initialize list temp_batchtemp\_batch
11:   while All processes have at least $b$ sequences based on all_q_listall\_q\_list do
12:     while temp_batchtemp\_batch does not have at least $b$ sequences do
13:      $temp\_batch\leftarrow temp\_batch.append(Q.dequeue())$
14:      q.deque()
15:     end while
16:     compute the the maximum per_device global batch size $b_{max}$ currently in all processes using all_q_listall\_q\_list (inferred, no collective operation)
17:     Pad temp_batchtemp\_batch to size $b_{max}$ by appending “pad sequences” as needed.
18:     Update all entries in all_q_listall\_q\_list based on inferred usage
19:     Perform one SSFT training step, SSFT ($\pi_{\bm{\theta}},temp\_batch,b_{max}$)
20:   end while
21:  end for
22:end for
Algorithm 1 presents the core SSFT implementation with optimal bipartite matching. In practice, the nested-loop computation used to populate $\mathrm{C}$ is fully vectorized and can be executed in a single forward pass. This does not blow up VRAM because (i) we do not store activations for these cost evaluations (no backprop through matching costs), and (ii) We only need to use the first ${\mathrm{T}_{\mathrm{L}}}<{\mathrm{T}_{{\mathbf{r}}}}$ NTP losses to compute the matching cost, as the NTP loss over the first few thousand tokens can already differentiate many reasoning traces in terms of their modes. Nevertheless, our code also supports matching over the full $\mathrm{T}_{{\mathbf{r}}}$} tokens by chunking the computation into a few batches, so this step does not become a VRAM bottleneck. Fine-tuning on 1k questions with 4 traces each, SSFT (optimal matching) took 6.5 h for 6 epochs, compared to 6.1 h for standard SFT, adding only $6.6\%$ overhead.
The primary VRAM bottleneck in SSFT remains the backpropagation Step 10, regardless of whether we use optimal or random matching, because the effective batch size scales with $\mathrm{M}$. To mitigate this, we split the backward pass into several gradient-accumulation steps. Although our experiments use the same number of reasoning traces per question, we also support variable number of targets using our queue-based batching in Algorithm 2. The complication arises when using distributed training with a variable-sized batch, as different processes require the same per-device batch size to perform collective operations. This is mitigated by padding with “PAD sequences” to align batch sizes. Our implementation minimizes the number of “PAD sequences” by storing a variable number of targets in a queue and dequeuing multiple items to form a per-device global batch, so smaller batches can be stitched together instead of always being padded.
A.2 Learned Matchings between Global Forking Tokens and Traces
Initially, any of the bipartite matching configuration $\bm{\sigma}_{k}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ can be computed as optimal, as the reserved global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ have no specific correlations with these traces $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. This can be observed in Figures 7 and 8 that $c(\bm{\sigma}_{k})$, the count of configuration $\bm{\sigma}_{k}$ being optimal during training, uniformly increases on all configuration indexes. However, as training goes with SSFT, we notice only a subset of $\bm{\mathfrak{S}}_{\mathrm{P}}$ accumulates mass. This indicates there are some unique correlations learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. This subset is denoted as $\bm{\mathfrak{S}}_{p}=\{\bm{\sigma}_{k}\}_{k=1}^{p}$, and we call the unique edges in $\bm{\mathfrak{S}}_{p}$ as learned matchings.
A.2.1 How to Choose the Global Forking Token for Pass@1
To find $\bm{\mathfrak{S}}_{p}$, we can simply track which configurations $\bm{\sigma}$ still accumulate mass in the last epoch. Then we can connect all the unique edges in $\bm{\mathfrak{S}}_{p}$ to visualize learned matchings. However, multiple global forking tokens may share the maximum number of connected edges in the learned matchings. To break the tie, we treat the counts as edge weights and select ${\textnormal{g}}^{(i)}$ with the largest weighted degree. We provide this implementation in our code.
A.3 Training details
A.3.1 Training Datasets
32B experiments with questions from s1(Main): We explain the process of generating our training dataset for experiments in Table 1, Figure 4, Figure 5, Figure 6, Figure 7. First, we use the 1000 questions from s1 (Muennighoff et al., 2025) and populate a pool of reasoning traces by distilling from GPT-OSS-120B-high reasoning, GPT-OSS-120B-medium reasoning, DeepSeek R1, Gemini Flash2.0 Thinking, and Claude4/4.1. We use temperature 1.0, maximum length of 32768, and sets high reasoning effort unless specified. We generate two traces per teacher model. We use Claude3.5 Sonnet to extract the answer from the distilled solutions and compare with the ground-truth answer. The correctness of these distilled traces are shown in 3.
For s1k-4mixed-reasoning dataset, we sample 4 traces per question from this pool, so the dataset consists of 1000 questions, each paired with 4 reasoning traces. This dataset was used to fine-tune SSFT-32B, SSFT-32B (random $\bm{\sigma}$), and SFT-mixed-distill-32B.
For fine-tuning SFT-OSS-distill-32B model, we only use the 1000 traces from “Run1” GPT-OSS-120B with high reasoning effort.
For obtaining the visualizations in Figure 7 and Figure 8, we fine-tune models using the same teacher models for all 1000 questions. we always choose the 4 traces from “Run1” of GPT-OSS-120B-high, GPT-OSS-120B-medium, DeepSeek R1, and Gemini Flash2.0 Thinking. As mentioned in Section 3.3 and Figure 7, we fine-tune under 3 bipartite matching hyperparameters to conduct this case study.
GPT-OSS-120B-high GPT-OSS-120B-medium DeepSeekR1 Gemini Flash2.0 Thinking Claude Opus4/4.1
Run1 796/1000 769/1000 620/1000 538/1000 656/1000
Run2 785/1000 753/1000 641/1000 545/1000 647/1000
Table 3: The number of correct reasoning traces distilled for the 1,000 questions in s1 by different teacher models. This evaluation is done by Sonnet comparing the predictions and the ground-truth answers. We see that GPT-OSS has the highest accuracy for s1 dataset.
A.3.2 Training Hyperparameters
For consistency, we use Qwen2.5-32B-Instruct (Yang et al., 2025a) as the base model for all of our 32B experiments. We use standard fine-tuning hyperparameters: we train for 6 epochs with a global batch size of 32, which is derived from 4 gradient accumulation steps and distributed training with 8 GPUS ($4\times 8=32$). This results in 756 gradient steps. The maximum sequence length is set to 32,768. We train with bfloat16 with a learning rate of 1e51e-5 warmed up linearly for 5% and then decayed to 0 using a cosine schedule. We choose AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay 1e41e-4. We only backpropagates the completion loss, which is the loss on reasoning traces and the answers. Fine-tuning SSFT-32B plus loggings took 6.5 hours on 8 NVIDIA B200 GPUs using PyTorch FSDP, Liger Kernel (Hsu et al., 2024) for fused cross entropy loss, and FlashAttention-2 (Dao, 2023) for fused attention computation. Fine-tuning SSFT-32B (random $\bm{\sigma}$) took 6.3 hours, and Fine-tuning SFT-mixed-distill-32B took 6.1 hours. Even our baseline SFT-OSS-distill-32B with only one trace per question, and our attempt to reproduce s1.1 took 1.66 hours, which is longer than the time reported by Muennighoff et al. (2025). This is due to using 8 GPUs instead of 16 GPUs, hardware and package differences. When training with s1k-4mixed-reasoning, we added one extra epoch from 5 epochs to 6 epochs, since we have 4x reasoning traces, but we did not linearly increase the number of epochs, as these traces can be similar, and the number of distinct questions is still 1,000. Overall, we made sure all of our models are fine-tuned with consistent hyperapameters.
A.3.3 Visualization of SSFT Training Dynamics
Figure 9 shows the standard training dynamics of SSFT with optimal bipartite matching. The resulting model is SSFT-32B. The loss plotted here is Equation 5.
Figure 9: Training dynamics of SSFT-32B on s1k-4mixed-reasoning
Figure 10 shows the evolution of bipartite matching during SSFT. Figures 10(a) and 10(b) show that the gap between optimal bipartite matching cost and non-optimal bipartite matching cost under other $\bm{\sigma}$ keeps widening during training. This means that these reasoning traces are indeed starting to match unique global forking tokens. Even though SSFT effectively optimizes a non-stationary objective which depends on model parameters $\bm{\theta}$}, the widening gap shows the inner discrete optimization is converging as training goes. Figure 10(c) also confirms that the model learned some unique correlations between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$, as only a subset of matchings are still computed as optimal. Compared to Figure 7, we see more $\bm{\sigma}$ accumulating mass towards the end. This is due to having more mixed diverse reasoning traces, so the model learned more intricate associations between these global forking tokens and truly diverse reasoning traces.
(a) The gap between the optimal bipartite matching cost in Equation 3 and second smallest bipartite matching cost. The increase shows the global forking tokens start preferring a specific matching.
(b) The gap between the optimal matching cost in Equation 3 and the average of bipartite matching cost under other $\bm{\sigma}$. The increase shows the global forking tokens start preferring a specific matching.
(c) Cumulative counts of $\bm{\sigma}_{k}$ computed as optimal over training. We observe that only a subset $\bm{\mathfrak{S}}_{p}$ accumulates mass towards the end, indicating some unique correlations learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
Figure 10: Dynamics of bipartite matching during the fine-tuning of SSFT-32B on s1k-4mixed-reasoning.
A.3.4 Ablation Study Training details (Removing High Quality Small Dataset)
For this ablation study, we choose Open-R1-Math220k default split, which has 93,000 math questions and $2\sim 4$ traces. Since Qwen2.5-Math-7B is a widely fine-tuned model using this dataset, we also choose it as our base model. We train for 3 epochs using 8 A100 GPUs, which took around 4 days. Our hyperparameters are mostly consistent with the recommended hyperparameters by Face (2025). We fine-tune both SSFT-OpenR1-93k-7B and SFT-OpenR1-93k-7B with a maximum length of 32768, learning rate of 4.0e054.0e-05 warmed up linearly for $3\%$ and decayed to 0 following cosine schedule, 8 gradient accumulation steps. For our SSFT method, we reserve $\mathrm{N}=4$ global forking tokens, and use the first 1000 tokens for matching. Again, only the completion loss is used for optimizing the model parameters.
A.4 Examples of SSFT-32B Reasoning Traces initiated by Different Global Forking Tokens
We randomly picked a set of six parallel generations from those where SSFT-32B correctly answered a problem from AIME 2025 under all six global forking tokens. This is question 11.
Problem 11 (AIME25) A piecewise linear periodic function is defined by $f(x)=\begin{cases}x,&\text{if }x\in[-1,1),\\[2.0pt] 2-x,&\text{if }x\in[1,3),\end{cases}\qquad\text{and}\qquad f(x+4)=f(x)\ \text{ for all real }x.$ The graph of $f(x)$ has a sawtooth pattern. The parabola $x=34y^{2}$ intersects the graph of $f(x)$ at finitely many points. The sum of the $y$-coordinates of these intersection points can be expressed in the form $\frac{a+b\sqrt{c}}{d},$ where $a,b,c,d$ are positive integers, $\gcd(a,b,d)=1$, and $c$ is squarefree. Find $a+b+c+d$.
Here are the 6 parallel generations by SSFT-32B. We see distinct reasoning traces produced by prompting with a unique ${\textnormal{g}}^{(i)}\in\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$. The differences are highlighted in red comments.
<think1> — Periodization + Quadratic Families + Vieta/Telescoping <|im_start|>assistant
<|im_start|>
<think1> “write $x=t+4k$ with $t\in[-1,3)$” … … “Case 1 $(t\in[-1,1),\,y=t)$ … Case 2 $(t\in[1,3),\,y=2-t)$” … … $34t^{2}-t-4k=0$ and $t=\dfrac{1\pm\sqrt{1+544k}}{68}$ … … “let u=2tu=2-t” and $u=\dfrac{-1\pm\sqrt{273+544k}}{68}$ … … Thus $u=\dfrac{-1+5\sqrt{185}}{68}$. … … “telescopes to rational constant 9/349/34” … … ([Comment:] periodize $x=t+4k$, solve two quadratic families, use Vieta pair-sums; edge-period gives the irrational $y$) </think1> <|im_start|>answer
Answer: …. 259
<think2> — Nearest-Multiple-of-4 Reduction <|im_start|>assistant
<|im_start|>
<think2> Start from $y=f(34y^{2})$ … … write $34y^{2}=4k+r$ with $r\in[-1,3)$ (choose $k=\big\lfloor\tfrac{34y^{2}+1}{4}\big\rfloor$) … … Branch A ($r\in[-1,1)$): $y=r\Rightarrow 34y^{2}-y-4k=0$ … … Branch B ($r\in[1,3)$): $y=2-r\Rightarrow 34y^{2}+y-(2+4k)=0$ … … Vieta: pair-sums $\tfrac{1}{34}$ and $\tfrac{1}{34}$ cancel over full periods … … only the edge/partial period remains ($k=8$, descending branch) … … $y=\dfrac{-1+\sqrt{273+544\cdot 8}}{68}=\dfrac{-1+5\sqrt{185}}{68}$ … … ([Comment:] reduce to $r\in[-1,3)$ via nearest multiple of 4; full periods cancel by $\pm\frac{1}{34}$ pair-sums; only the boundary $k=8$ root contributes) </think2> <|im_start|>answer
Answer: …. 259
<think3> — Paper-Style Sections; Explicit Pair-Sums and Totals <|im_start|>assistant
<|im_start|>
<think3> “Now we need sum of all $y$-coordinates” … … $\sum_{n=0}^{8}y_{1,n}+\sum_{n=0}^{7}y_{C,n}$ … … $y=\dfrac{-1+\sqrt{4625}}{68}$ … … “Therefore $S_{2}=\dfrac{5\sqrt{185}-17}{68}$” … … “Total sum $S$ …” … … ([Comment:] formal write-up; separates sums by family, explicitly displays the boundary $y=\frac{-1+\sqrt{4625}}{68}$, then totals) </think3> <|im_start|>answer
Answer: …. 259
<think4> — Set/System Framing Before Algebra <|im_start|>assistant
<|im_start|>
<think4> “the set $\{(x,f(x))\}$ … Intersecting it with the parabola $x=34y^{2}$ requires solving $y=f(x),\ x=34y^{2}$.” … … (then proceeds to the same two-case algebra) … … $u_{+}=\dfrac{-1+\sqrt{273+544\cdot 8}}{68}=\dfrac{-1+5\sqrt{185}}{68}$. … … ([Comment:] definition-first framing; after setting up the system, it reaches the same boundary $u_{+}$ value on the partial period) </think4> <|im_start|>answer
Answer: …. 259
<think5> — Outline & Symmetry: Full Periods Cancel, Last Period Contributes <|im_start|>assistant
<|im_start|>
<think5> “the sum is zero … That’s a key insight.” … … “for each full period $(k=0$ to $k=7)$, the sum … is zero … now consider the partial period k=8k=8” … … Compute $u_{\text{pos}}(k{=}8)=\dfrac{-1+\sqrt{4625}}{68}$. … … (term-by-term cancellation examples for $k=0,1$) … … ([Comment:] symmetry explains full-period cancellation; the lone boundary contribution at $k=8$ supplies the irrational term) </think5> <|im_start|>answer
Answer: …. 259
<think6> — Stream-of-Consciousness Scratchwork <|im_start|>assistant
<|im_start|>
<think6> “Okay, so I need to solve this problem … Let me try to break this down step by step.” … … “Wait, actually, no. Wait,” … … $y_{A,8}=\dfrac{-1+\sqrt{4625}}{68}$. … … ([Comment:] exploratory self-checks; reaches the same boundary $y$ on the partial period) </think6> <|im_start|>answer
Answer: …. 259
A.5 Alternative Implementation of training with multi-target sequences
Figure 11 shows an alternative implementation for training with multiple target parallel sequences. Instead of flattening multi-target data along the batch dimension, we can also concatenate parallel reasoning traces along the sequence dimension, modify the causal attention matrix, and position ids for positional embeddings. While easy to implement, this version extends the sequence length and cannot perform gradient accumulation along the batch dimension to maintain the same VRAM as training with a single reasoning trace. Since most fine-tuning already sets the micro-batch size to 1 due to the increased number of reasoning tokens already present in a single reasoning path, this implementation also cannot further shrink the micro-batch size to accommodate the extended context length. In addition, the code that uses the initial flash-attention implementation, which requires causal attention matrix, is not compatible with this setup. In terms of choosing the number of sequences in a global batch size, this implementation is also less flexible.
Figure 11: An illustration of training with multiple parallel targets along the sequence dimension.